OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (–4, 0) and a directrix at x = 4

OpenStudy (anonymous):

@ChipperJay*

OpenStudy (anonymous):

That means the vertex is the midpoint of the directrix and the focus which means that it is (0,0), so that and the formula is y=(1/4n)x^2 and n is 4 because that is the distance from the directrix to the vertex. So the equation is y=(1/16)x^2.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

@rrrrr

OpenStudy (anonymous):

what about this one

OpenStudy (anonymous):

A building has an entry the shape of a parabolic arch 96 ft high and 18 ft wide at the base as shown below.

OpenStudy (anonymous):

Find an equation for the parabola if the vertex is put at the origin of the coordinate system.

OpenStudy (anonymous):

my answer was y = (-32/27)(x - 0)^2 + 0

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

I'm not sure for this one, you should ask someone else :( sorry

OpenStudy (anonymous):

oh :( alright thanks anyway!

OpenStudy (anonymous):

could I ask you another question similar to the first? :)

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

just wanna make sure I've got the right answers and my work is correct :)

OpenStudy (anonymous):

I don't know this question. I apologize for this, but thanks for mentioning me anyway. I feel special. :)

OpenStudy (anonymous):

sure :D

OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a vertex at the origin and a focus at (0, 9). and I got: y=1/36x^2

OpenStudy (anonymous):

aw, it's alright. thanks anyway @ChipperJay* :)

OpenStudy (anonymous):

yes because 1/4n and n=9 so 1/36 so that is correct

OpenStudy (anonymous):

yes! thanks so much. :)

OpenStudy (anonymous):

youre welcome

OpenStudy (anonymous):

sorry could I ask you just oneeee more final question?

OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, 8) and a directrix at y = -8

OpenStudy (anonymous):

@doppler @rrrrr @ChipperJay* :)

OpenStudy (anonymous):

ok that will be y=(1/32)x^2 using the formula I mentioned

OpenStudy (anonymous):

alright thanks :)